# The Landscape of GPU-Centric Communication

### v2

@@ 6. Conclusion
The discussion includes an analysis of vendor-provided communication and memory management mechanisms for multi-GPU execution, as well as a review of major communication libraries such as CUDA-aware MPI, NCCL/RCCL, NVSHMEM, and ROC_SHMEM, highlighting their advantages, challenges, and performance considerations. Furthermore, the paper presents important research paradigms such as CPU-free networking, debugging tools for communication, and discusses future directions, and unresolved questions. By thoroughly examining GPU-centric communication techniques across both software and hardware layers, this paper aims to equip researchers, developers, engineers, and library designers with the knowledge needed to fully leverage multi-GPU systems.
Acknowledgement
Authors at Koç University were supported from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 949587).

### v3

@@ 6. Conclusion
The discussion includes an analysis of vendor-provided communication and memory management mechanisms for multi-GPU execution, as well as a review of major communication libraries such as CUDA-aware MPI, NCCL/RCCL/oneCLL, NVSHMEM, ROC_SHMEM, and Intel SHMEM highlighting their advantages, challenges, and performance considerations. Furthermore, the paper presents important research paradigms such as CPU-free networking, debugging tools for communication, and discusses future directions, and unresolved questions. By thoroughly examining GPU-centric communication techniques across both software and hardware layers, this paper aims to equip researchers, developers, engineers, and library designers with the knowledge needed to fully leverage multi-GPU systems.
Acknowledgment
Authors at Koç University were supported from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 949587). Flavio Vella and Daniele De Sensi are supported by the European Union’s Horizon Europe under grant 101175702 (NET4EXA). Daniele De Sensi is supported by Sapienza University Grants ADAGIO and D2QNeT (Bando per la ricerca di Ateneo 2023 and 2024).